OpenStudy (anonymous):

Romo makes $12 an hour for his first hour of work, $17 on his second, $22 on his third hour an so on. How much money will he earn on his 50th hour of work?

OpenStudy (tkhunny):

Sift through it an build it! 1:12 = 12 + 0 2:17 = 12 + 5 3:22 = 12 + 10 = 12 + 2(5) 4:27 = 12 + 3(5) ... n:?? = 12 + (n-1)(5)